Menlo Company distributes a single product. The companys sales and expenses for last month follow: Total $ 318,000 222,600 PReq 1 1 Reqz Reg 2 | Req3A Req 3A Req 3B Req 38 Req 4 Rega Reg 5 Reqs Without resorting to computations, what is the total coReq 1 Reg 2 Req ЗА Req 3B Req 4 Req 5 How many units would have to be sold each month to attain a target profit of $38,400? UReq 1 Reg 2 Req 3A Req 3B Reg 4 Reg 5 Verify your answer by preparing a contribution format income statement at the target saReq 1 Reg 2 Req 3A Req 3B Req 4 Reg 5 Refer to the original data. Compute the companys margin of safety in both dollar and pReq 1 Reg 2 Req 3A Req 3B Reg 4 Reg 5 What is the companys CM ratio? If sales increase by $82,000 per month and there is no
Answer #1
Formula Calculation
Req-1 Break-even point in unit sales 12500 Units Fixed cost/Contribution per unit 75,000/6
Break-even point in Dollar Sales $250,000 Fixed Cost/Contribution ratio 250,000/30%
Contribution ratio = 6/20 = 30%
Req-2 Total contribution margin at break-even $75,000 At break-even total contribution margin = fixed Cost
Req-3A Units sales needed to attain target profit                        18,900 (Target profit+fixed cost)/contribution per unit (38,400+75,000)/6
Req-3B Menio Company
Contribution Income Statement
Total Per Unit
Sales $378,000 20 (18,900*20)
Variable expense $264,600 14 (18,900*14)
Contribution margin $113,400 6
Fixed expense $75,000
$38,400
Dollars Percentage
Req-4 Margin of Safety $68,000 21.38%
(318,000-250,000) Margin of safety =   (Current sales level – breakeven point) / Current sales level X 100
Req-5 CM Ratio 30% (Sales-Variable expense)/sales
Net operating income increased by $24,600 (82,000*30%)

*if you need any further help in understanding any of the calculation above please feel free to ask in the comment section. Please give your valuable feedback.